# LIST: Learning to Index Spatio-Textual Data for Embedding based Spatial Keyword Queries

### 12 Mar 2024 v1

@@ 18       $Res_{q}\leftarrow\text{argTop-k}_{o\in C_{i}}ST(q,o)$;
For a given query, the time complexity of LIST is $O((l-2)d^{2}+dc+\frac{n}{c}(d+2)+d)$. $O((l-2)d^{2}+dc)$ is the time complexity for our index (lines 15), which is the inference computation cost of Equation 11. $O(\frac{n}{c}(d+2)+d)$ is the time cost of our relevance model (line 17). $\frac{n}{c}$ denotes the number of objects to be checked by the relevance model, which is approximately $\frac{1}{c}$ of the entire dataset. This is because of the even cluster distribution (Verified in Section V-E), which means approximately $\frac{1}{c}$ of the dataset needs to be visited. Specifically, $O(\frac{nd}{c})$ is the computational cost of Equation 3. $O(\frac{2n}{c})$ is the computational cost of Equation 5, and $O(d)$ is the computational cost of Equation 6.
@@
V EXPERIMENTS
TABLE I: Datasets Statistics.
@@ V EXPERIMENTS
  • RQ3: How does LIST perform on datasets and training datasets of varying sizes?
@@
TABLE II: Comparison of relevance models across three datasets by brute-force search.
@@
V-A Experimental Setup
Datasets. We utilize three benchmark datasets: Beijing, Shanghai, and Geo-Glue. The Beijing and Shanghai datasets [5] are provided by Meituan, a Chinese retail services platform. Users submit queries consisting of location and keywords. The clicked Points of Interest (POIs) recorded in the query log are regarded as the ground truth answers to these queries.Note that the click-through data recorded in the search log is widely used as ground-truth for information retrieval evaluation as explicit feedback such as ratings is challenging to collect and implicit feedback such as click-through data can be tracked automatically and is thus much easier to collect for content providers [64]. Both datasets are anonymized to protect user privacy and more details can be found here [5]. The Geo-Glue dataset [65, 12] is released by DAMO Academic. The POIs in the dataset are crawled from OpenStreetMap [^1] https://www.openstreetmap.org/, and the queries and the corresponding ground truth POIs are manually generated by domain experts. Notably, in the released Geo-Glue dataset, the coordinates of objects and queries are modified due to privacy considerations, which results in many objects with identical geo-locations.
Note that, due to privacy considerations, the release of query log datasets from the industry is restricted. As a result, to our knowledge, no other datasets containing ground-truth query results have been released, except for the three datasets used in our experiment. To investigate the scalability of the proposed framework, we conduct a scalability study to show LIST’s efficiency on larger datasets, where the Geo-Glue dataset is augmented with more crawled POIs (shown in Section V-D)
Dataset Split. The statistics of datasets are stated in Table I, where each record represents a single ground-truth label between an object and a query, and each query may have multiple ground-truth objects. For the Beijing and Shanghai datasets, to ensure a fair comparison, we follow the previous split strategy [5], where 90% of queries and their corresponding ground-truth records are used as the training set and the remaining queries as the test set. From the training set, we randomly choose 10% data as the validation set to tune hyperparameters. For the Geo-Glue dataset [12], we follow the provided splits for training, validation, and test data.
Effectiveness metric. Following previous studies [12, 5], we use two metrics, i.e., the Recall and Normalized Discounted Cumulative Gain (NDCG), to evaluate the effectiveness. Recall@$k$ evaluates the proportion of positive objects contained in the top-$k$ candidates for a given query. NDCG@$k$ considers the order of ground-truth objects in the retrieved objects, reflecting the quality of the ranking in the retrieved list. We assign the graded relevance of the result at position $i$ as $rel_{i}\in\{0,1\}$, where $rel_{i}=1$ when the object is relevant to the query, otherwise $rel_{i}=0$. More details can be found in [5]. Specifically, the following metrics are utilized: Recall@10, Recall@20, NDCG@1, and NDCG@5.
Baselines of Relevance Model. We compare our proposed relevance model, denoted as LIST-R, with various state-of-the-art spatio-textual relevance models or state-of-the-art deep textual relevance models. We also modify the state-of-the-art textual relevance model DPR [38] into ${}_{D}$ to answer TkQs.
@@ V-A Experimental Setup
  • PALM [13]: This method employs deep neural networks for query-object spatio-textual relevance.
  • DrW [5]: This is the newest deep relevance based method for answering TkQs.
  • DPR [38]: This is a state-of-the-art textual relevance model for document retrieval. This method does not consider spatial relevance.
@@
  • MGeo [12]: This is also the newest deep relevance based method for answering TkQs. Note that we cannot reproduce the experimental results reported in the paper, and therefore we use the evaluation results from the original paper.
Baselines of LIST. We compare LIST with the following retrieve and rerank methods. We select the IR-Tree [2] to accelerate the TkQ search, serving as an index baseline. There exist two categories of ANNS indexes, as discussed in Section II, and our method belongs to the first category. We select the state-of-the-art ANNS indexes in the first category, i.e., IVF [46], LSH [63], and HNSW [52], as index baselines. For the second category of methods, such as product quantization based methods, their standalone efficiency is no better than a brute-force search [^2] https://github.com/facebookresearch/faiss/issues/148. Consequently, following previous work [46], we choose IVFPQ, which integrates PQ with the IVF index from the first category, which is considered as a state-of-the-art solution. Additionally, existing ANNS indexes consider only textual embedding but not the spatial factor. To integrate the spatial factor, we use ${}_{D}$, which is the most effective baseline of the spatio-textual relevance model, to rerank the objects retrieved by an index.
An important hyperparameter is $k$, which represents the number of objects retrieved by the indexes and reranks by the ${}_{D}$. Specifically, on the Beijing and Shanghai datasets, for a fair comparison, we set $k$ to be 10,000, which is the average number of objects retrieved by our index per query. On the geo-glue dataset, $k$ is set to be 55,000 for the same reason above. We evaluate the effect of $k$ in Section V-C.
@@
  • LSH+${}_{D}$: This method constructs an LSH index with the embeddings produced by the ${}_{D}$, which retrieves the top-$k$ relevant objects by fetching similar embeddings in the same buckets, and reranks the retrieved objects by the ${}_{D}$. Following previous work [63], we set the length of the hash code $nbits$ to 128.
  • HNSW+${}_{D}$: This method constructs an HNSW index over the embeddings produced by the ${}_{D}$, which retrieves the top-$k$ relevant objects by conducting beam searches over the proximity graph, and reranks the retrieved objects by the ${}_{D}$. Following previous work [52], we set the number of links $M$ to 48 and $efConstruction$ to 100.
  • IVFPQ+${}_{D}$: This method integrates the IVF index with the product quantization technique [46] to retrieve objects. The retrieved objects are reranked by the ${}_{D}$. We set the number of clusters $c$ to be the same as our index. Following the instruction [63], we set the number of centroids $w$ to 32 and the number of bits $nbits$ to 8, the number of clusters to search $cr$ at 4. Here, we do not set $cr$ the same as our index since the product quantization accelerates computations within clusters, allowing access to more clusters with comparable time costs. Notably, the maximal $k$ supported by the Faiss library is 2,048, therefore we set $k$ to 2,048 across three datasets.
Due to the page limitation, we demonstrate the effect of key parameters for all these indexes in our full version paper [^3] https://github.com/Heisenberg-Yin/LIST.
Implementations. The relevance model and the index are trained using Pytorch. During the query phase, the index and all relevance models are inferred in C++ by the ONNX system [^4] https://github.com/onnx/onnx. In our model, we utilize the bert-base-Chinese pre-trained model [^5] https://huggingface.co/bert-base-chinese from the huggingface Library [66] as encoders, which is the same as previous work [5, 12]. The footstep controller $t$ of $\mathcal{T}$ is set to be 1000 for all three datasets. $\text{neg}_{\text{start}}$ is set to 50,000 for the Beijing dataset, 60,000 for the Shanghai dataset, and 200,000 for the Geo-Glue dataset. We empirically set the cluster number $c$ to approximately $\frac{n}{10,000}$, i.e., 20 for the Beijing and Shanghai datasets and 300 for the Geo-Glue dataset. This is because ranking 10,000 objects by the relevance model is computationally feasible and does not notably compromise effectiveness. The number of clusters to route $cr$ for queries and objects is set to 1 by default, with different $cr$ and $c$ settings shown in Section V-C. The implementations of DrW are from publicly available source codes, and we make use of the implementation of IVF, LSH, IVFPQ, and HNSW provided by the Faiss library [63], while others are implemented by ourselves. The Faiss library is implemented in C++, providing a fair comparison with our index. Our default experiment environment is CPU: Intel(R) Core(TM) i9-10900X CPU@3.70GHz, Memory: 128G, and GPU: RTX3080 10GB.
TABLE III: Comparison of retrievers across three datasets.
@@
V-B Effectiveness Study (RQ1)
Effectiveness of Relevance Model LIST-R. To validate our proposed relevance model’s effectiveness, denoted as LIST-R, we compare it with other relevance models across the three datasets. All relevance models perform a brute-force search over the entire dataset to identify the top-k objects. Table II reports the effectiveness of the evaluated methods. DrW and PLAM cannot be evaluated on the Geo-Glue dataset via brute-force search because of their slow querying speed, requiring more than a day for evaluation. We have the following findings: (1) LIST-R consistently outperforms all the baseline models on all three datasets across every metric. Specifically, LIST-R achieves up to a 19.21% improvement over the best baseline on NDCG@1 and up to a 12.79% improvement on recall@10. (2) Traditional ranking methods are less effective than deep relevance models. On the three datasets, TkQ is outperformed by DrW and ${}_{D}$. This can be attributed to the word mismatch issue discussed in Section III-B. Notably, PALM is outperformed by the TkQ, which is consistent with the results reported by [5]. (3) Directly applying the dual-encoder model to our task is less effective. On the Beijing and Shanghai datasets, DPR is outperformed by TkQ, which is caused by the ignorance of spatial relevance. DPR’s enhanced effectiveness on the geo-glue dataset can be attributed to the property of the dataset, where most queries are address descriptions [12].
Effectiveness of LIST. We investigate the effectiveness of LIST by comparing it with the state-of-the-art baselines. The evaluation results are reported in Table III. We have the following findings: (1) Directly applying ANNS indexes in TkQ task results in severe effectiveness degradation. These inferior results can be attributed to the ignorance of the spatial factor by these ANNS indexes. Compared with other ANNS indexes-based methods, TkQ+${}_{D}$ offers better effectiveness across the three datasets. (2) Overall, LIST significantly outperforms all the baselines, achieving the highest scores on all datasets. Compared with the strongest baseline TkQ+${}_{D}$, LIST achieves up to 12.13% improvements on Recall@10 metric and 19.20% on NDCG@1 metric.
TABLE IV: Query runtime on three datasets (ms).
@@ V-B Effectiveness Study (RQ1)
V-C Efficiency Study (RQ2)
Runtime of LIST. We investigate the querying speed of LIST by comparing it with the state-of-the-art baselines. During the query phase, our proposed index, IVF, LSH, IVFPQ, and HNSW remain in the GPU, while IR-Tree remains in the memory. The total runtime is computed by calculating the average time cost of 5,000 randomly sampled queries from the test query set. We present the total runtime of different methods in Table IV. Based on the experimental results, we have the following findings: (1) Compared with the most effective baselines, i.e., TkQ+${}_{D}$, LIST is one order of magnitude faster on the Beijing and Shanghai datasets and three orders of magnitude faster on the Geo-Glue dataset. (2) IVF+${}_{D}$ and IVFPQ+${}_{D}$ have similar runtimes as LIST. However, their effectiveness is much worse than LIST as stated in Table III. Specifically, compared with the IVFPQ+${}_{D}$, LIST achieves an improvement of up to 22.70% on the NDCG@1 metric across the three datasets. Compared with the IVF+${}_{D}$ baseline, we observe an improvement of up to 77.61% on the NDCG@1 metric across the three datasets. (3) The TkQ+${}_{D}$ exhibits a significantly slower querying speed than TkQ across three datasets. The reason is that IR-Tree exploits a priority queue, using the bottom object to prune lower-scoring objects. As the queue size (hyperparameter $k$) increases, the IR-Tree’s pruning ability declines, aligning with previous empirical results [2].
@@ V-C Efficiency Study (RQ2)
Trade-off Study by varying $k$ and $c$. We examine LIST’s ability to trade off effectiveness and efficiency under the different number of cluster $k$ settings. Specifically, on the Beijing and Shanghai datasets, we set the hyperparameter cluster number $c$ of LIST, IVF+${}_{D}$, and IVFPQ+${}_{D}$ from 5, 10, 20, 50, to 100. For other baselines, i.e., LSH+${}_{D}$, HNSW+${}_{D}$, and TkQ+${}_{D}$, we set hyperparameter top-$k$ from 40,000, 20,000, 1,$0$, 5,000 to 2,000. This ensures the number of objects retrieved for reranking is consistent across different baselines, leading to a fair comparison. Other hyperparameter settings are fixed. We present the trade-off results on the NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 4 while other metrics and datasets have similar results. We have the following findings: (1) LIST outperforms other methods on both effectiveness and efficiency. This demonstrates LIST’s capability of maintaining effectiveness while accelerating querying speed. (2) The effectiveness of TkQ+${}_{D}$ is consistent as $k$ increases while time overhead increases significantly. This is due to the word mismatch issue discussed in Section III-B. Many relevant objects without word overlap with the query remain hard to retrieve as $k$ increases. (3) HNSW+${}_{D}$ exhibits almost constant results. This is because HNSW builds a sparse proximity graph, limiting the reachable objects for each query. Therefore, when $k$ exceeds the number of reachable objects, its performance remains unchanged.
@@
Trade-off Study by varying $k$ and $cr$. We further investigate the impact of the number of clusters $cr$ to route for queries and objects. We vary $cr$ from 1, 2, 3, to 4 for IVF+${}_{D}$ and our index, from 4, 5, 6, to 7 for IVFPQ+${}_{D}$. For other baselines, we set hyperparameter top-$k$ from 10,000, 20,000, 3,$0$, to 4,$0$ for a fair comparison like the previous trade-off study. Other hyperparameter settings are fixed. Similarly, we report the trade-off results on NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 5. According to the experiment results, LIST outperforms other methods in both effectiveness and efficiency under different $cr$ settings. Increasing $cr$ can enhance effectiveness, but it also leads to higher latency. In practice, users can adjust this parameter according to their needs.
Memory Consumption of LIST. The memory consumption of LIST is composed of three parts: the memory used by the proposed relevance model, the memory used by the proposed index, and the memory used for object text embeddings that are produced in advance. The experiment results are presented in Table V, which demonstrates the remarkable memory efficiency of LIST. Compared with LSH+${}_{D}$, TkQ+${}_{D}$, and HNSW+${}_{D}$, LIST requires less memory. The reason is that our index stores only a lightweight MLP $c$-cluster classifier. This storage requirement is less than the memory consumption of the proximity graph of HNSW, hash tables of LSH, and the inverted file of the IR-Tree.
TABLE V: Memory usage on three datasets (MB).
@@
V-D Effect of Dataset Size (RQ3)
Scalability Study. We further explore the scalability of LIST and the proposed relevance model LIST-R. We supplement the Geo-Glue dataset with new POIs crawled from Open Street Maps in Hangzhou. For the newly acquired POIs, we utilize the trained dual-encoder module and index to partition them into distinct clusters. We report the runtime of LIST and LIST-R on Figure 6. We observe that, as the number of objects increases, the runtime of LIST and LIST-R scales linearly.

### 18 Mar 2024 v2

@@ 18       $Res_{q}\leftarrow\text{argTop-k}_{o\in C_{i}}ST(q,o)$;
For a given query, the time complexity of LIST is $O((l-2)d^{2}+dc+\frac{n}{c}(d+2)+d)$. $O((l-2)d^{2}+dc)$ is the time complexity for our index (lines 15), which is the inference computation cost of Equation 11. $O(\frac{n}{c}(d+2)+d)$ is the time cost of our relevance model (line 17). $\frac{n}{c}$ denotes the number of objects to be checked by the relevance model, which is approximately $\frac{1}{c}$ of the entire dataset. This is because of the even cluster distribution (Verified in Section 5.5), which means approximately $\frac{1}{c}$ of the dataset needs to be visited. Specifically, $O(\frac{nd}{c})$ is the computational cost of Equation 3. $O(\frac{2n}{c})$ is the computational cost of Equation 5, and $O(d)$ is the computational cost of Equation 6.
@@
5 Experiments
Table 1: Datasets Statistics.
@@ 5 Experiments
  • RQ3: How does LIST scale with the size of dataset?
@@
5.1 Experimental Setup
Datasets. To evaluate the effectiveness and efficiency of our proposed retriever LIST, we utilize three benchmark datasets: Beijing, Shanghai, and Geo-Glue. The Beijing and Shanghai datasets liu2023effectiveness are provided by Meituan, a Chinese retail services platform. Users submit a query through the platform, which consists of a query location and a set of keywords. Subsequently, the Points of Interest (POIs) that users clicked on are recorded in the query log, and is considered as a ground-truth. Note that the click-through data recorded in the search log may be the only feasible way to get a large scale of ground truth data for spatial keyword queries liu2023effectiveness . The explicit feedback such as ratings is very challenging to collect DBLP:conf/www/HeLZNHC17 . Note that using query logs is also the popular way of generating ground truth in the Information Retrieval literature DBLP:conf/www/HeLZNHC17 ; DBLP:conf/cikm/HuangHGDAH13 ; DBLP:conf/www/YaoTCYXD021 ; DBLP:conf/kdd/Joachims02 ; DBLP:journals/csur/WuSZXC23 . Therefore, following the previous work liu2023effectiveness , we utilize the two datasets to evaluate our proposed solution LIST and treat the clicked POIs as ground truth relevant objects to the corresponding queries. The Geo-Glue dataset li2023geoglue ; DBLP:conf/sigir/DingCXHLZX23 is released by DAMO Academic. The POIs in the Geo-Glue dataset are crawled from OpenStreetMap [^1] https://www.openstreetmap.org/, and the queries and the corresponding ground truth POIs are manually generated by domain experts. Notably, in the released Geo-Glue dataset, the coordinates of objects and queries are modified due to privacy considerations, which results in many objects with identical geo-locations.
The release of query log datasets from the industry is highly restricted. As a result, to the best of our knowledge, there are no other public datasets that contain ground-truth query results or query logs, except for the three datasets used in our experiments. To investigate the scalability of the proposed framework, we conduct a scalability study to show LIST’s efficiency on larger datasets, where the Geo-Glue dataset is augmented with more crawled POIs (shown in Section 5.4)
Dataset Split. The statistics of datasets are stated in Table 1, where each record represents a single ground-truth label between an object and a query, and each query may have multiple ground-truth objects. For the Beijing and Shanghai datasets, to ensure a fair comparison, we follow the previous split strategy liu2023effectiveness , where 90% of queries and their corresponding ground-truth records are used as the training set and the remaining queries as the test set. From the training set, we randomly choose 10% data as the validation set to tune hyperparameters. For the Geo-Glue dataset DBLP:conf/sigir/DingCXHLZX23 , we follow the provided splits for training, validation, and test data.
Table 2: Comparison of relevance models across three datasets by brute-force search.
@@ 5.1 Experimental Setup
Effectiveness metric. Following previous studies DBLP:conf/sigir/DingCXHLZX23 ; liu2023effectiveness , we use two metrics, i.e., the Recall and Normalized Discounted Cumulative Gain (NDCG), to evaluate the effectiveness. Recall@$k$ evaluates the proportion of positive objects contained in the top-$k$ candidates for a given query. NDCG@$k$ considers the order of ground-truth objects in the retrieved objects, reflecting the quality of the ranking in the retrieved list. We assign the graded relevance of the result at position $i$ as $rel_{i}\in\{0,1\}$, where $rel_{i}=1$ when the object is relevant to the query, otherwise $rel_{i}=0$. More details can be found in liu2023effectiveness . Specifically, the following metrics are utilized: Recall@10, Recall@20, NDCG@1, and NDCG@5.
Baselines of Relevance Model. We compare our proposed relevance model, denoted as LIST-R, with various state-of-the-art spatio-textual relevance models or state-of-the-art deep textual relevance models. We also modify the state-of-the-art textual relevance model DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 into ${}_{D}$ to answer TkQs.
@@
  • PALM DBLP:conf/aaai/ZhaoPWCYZMCYQ19 : This method employs deep neural networks for query-object spatio-textual relevance.
  • DrW liu2023effectiveness : This is the newest deep relevance based method for answering TkQs.
  • DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 : This is a state-of-the-art textual relevance model for document retrieval. This method does not consider spatial relevance.
  • ${}_{D}$: This method replaces the text relevance computation in DrW liu2023effectiveness with DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 .
  • MGeo DBLP:conf/sigir/DingCXHLZX23 : This is a recent deep relevance based method for answering TkQs. Note that we cannot reproduce the experimental results reported in the paper by running the official code, and we use the evaluation results from the original paper.
@@
An important hyperparameter is $k$, which represents the number of objects retrieved by the indexes and reranks by ${}_{D}$. Specifically, on the Beijing and Shanghai datasets, for a fair comparison, we set $k$ to be 10,000, which is the average number of objects retrieved by our index per query. On the geo-glue dataset, $k$ is set to be 55,000 for the same reason above. We evaluate the effect of $k$ in Section 5.3.
@@
[…]
  • LSH+${}_{D}$: This method constructs an LSH index with the embeddings produced by ${}_{D}$, which retrieves the top-$k$ relevant objects by fetching similar embeddings in the same buckets, and reranks the retrieved objects by ${}_{D}$. Following previous work johnson2019billion , we set the length of the hash code $nbits$ to 128.
  • HNSW+${}_{D}$: This method constructs an HNSW index over the embeddings produced by ${}_{D}$, which retrieves the top-$k$ relevant objects by conducting beam searches over the proximity graph, and reranks the retrieved objects by ${}_{D}$. Following previous work malkovEfficientRobustApproximate2020 , we set the number of links $M$ to 48 and $efConstruction$ to 100.
  • IVFPQ+${}_{D}$: This method integrates the IVF index with the product quantization technique jegouProductQuantizationNearest2011 to retrieve objects. The retrieved objects are reranked by ${}_{D}$. We set the number of clusters $c$ to be the same as our index. Following the instruction johnson2019billion , we set the number of centroids $w$ to 32 and the number of bits $nbits$ to 8, the number of clusters to search $cr$ at 4. Here, we set a larger value for $cr$ than that in our index since the product quantization accelerates computations within clusters, allowing access to more clusters with comparable time costs. Notably, the maximal $k$ supported by the Faiss library is 2,048, therefore we set $k$ to 2,048 across three datasets.
Table 3: Comparison of retrievers across three datasets.
@@
Implementations. The relevance model and the index are trained using Pytorch. During the query phase, the index and all relevance models are inferred in C++ by the ONNX system [^3] https://github.com/onnx/onnx. In our model, we utilize the bert-base-Chinese pre-trained model [^4] https://huggingface.co/bert-base-chinese from the huggingface Library wolf2019huggingface as encoders, which is the same as previous work liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 . The footstep controller $t$ of $\mathcal{T}$ is set to be 1000 for all three datasets. $\text{neg}_{\text{start}}$ is set to 50,000 for the Beijing dataset, 60,000 for the Shanghai dataset, and 200,000 for the Geo-Glue dataset. We empirically set the cluster number $c$ to approximately $\frac{n}{10,000}$, i.e., 20 for the Beijing and Shanghai datasets and 300 for the Geo-Glue dataset. This is because ranking 10,000 objects by the relevance model is computationally feasible and does not notably compromise effectiveness. The number of clusters to route $cr$ for queries and objects is set to 1 by default, with different $cr$ and $c$ settings shown in Section 5.3. The implementations of DrW are from publicly available source codes, and we make use of the implementation of IVF, LSH, IVFPQ, and HNSW provided by the Faiss library johnson2019billion , while others are implemented by ourselves. The Faiss library is implemented in C++, providing a fair comparison with our index. Our default experiment environment is CPU: Intel(R) Core(TM) i9-10900X CPU@3.70GHz, Memory: 128G, and GPU: RTX3080 10GB.
5.2 Effectiveness Study (RQ1)
Effectiveness of Relevance Model LIST-R. To validate our proposed relevance model’s effectiveness, denoted as LIST-R, we compare it with other relevance models across the three datasets. All relevance models perform a brute-force search over the entire dataset to identify the top-k objects. Table 2 reports the effectiveness of the evaluated methods. DrW and PLAM cannot be evaluated on the Geo-Glue dataset via brute-force search because of their slow querying speed, requiring more than a day for evaluation. We have the following findings: (1) LIST-R consistently outperforms all the baseline models on all three datasets across every metric. Specifically, LIST-R achieves up to a 19.21% improvement over the best baseline on NDCG@1 and up to a 12.79% improvement on recall@10. (2) Traditional ranking methods are less effective than deep relevance models. On the three datasets, TkQ is outperformed by DrW and ${}_{D}$. This can be attributed to the word mismatch issue discussed in Section 3.2. Notably, PALM is outperformed by the TkQ, which is consistent with the results reported by liu2023effectiveness . (3) Directly applying the dual-encoder model to our task is less effective. On the Beijing and Shanghai datasets, DPR is outperformed by TkQ, which is caused by the ignorance of spatial relevance. DPR’s enhanced effectiveness on the geo-glue dataset can be attributed to the property of the dataset, where most queries are address descriptions DBLP:conf/sigir/DingCXHLZX23 .
[…]
Table 4: Query runtime on three datasets (ms).
@@ 5.2 Effectiveness Study (RQ1)
5.3 Efficiency Study (RQ2)
Runtime of LIST. We investigate the querying speed of LIST by comparing it with the state-of-the-art baselines. During the query phase, our proposed index, IVF, LSH, IVFPQ, and HNSW remain in the GPU, while IR-Tree remains in the memory. The total runtime is computed by calculating the average time cost of 5,000 randomly sampled queries from the test query set. We present the total runtime of different methods in Table 4. Based on the experimental results, we have the following findings: (1) Compared with the most effective baselines, i.e., TkQ+${}_{D}$, LIST is one order of magnitude faster on the Beijing and Shanghai datasets and three orders of magnitude faster on the Geo-Glue dataset. (2) IVF+${}_{D}$ and IVFPQ+${}_{D}$ have similar runtimes as LIST. However, their effectiveness is much worse than LIST as stated in Table 3. Specifically, compared with the IVFPQ+${}_{D}$, LIST achieves an improvement of up to 22.70% on the NDCG@1 metric across the three datasets. Compared with the IVF+${}_{D}$ baseline, we observe an improvement of up to 77.61% on the NDCG@1 metric across the three datasets.
@@ 5.3 Efficiency Study (RQ2)
Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters ($c$). We examine LIST’s ability to trade off effectiveness and efficiency under the different number of cluster $k$ settings. Specifically, on the Beijing and Shanghai datasets, we set the hyperparameter cluster number $c$ of LIST, IVF+${}_{D}$, and IVFPQ+${}_{D}$ from 5, 10, 20, 50, to 100. For other baselines, i.e., LSH+${}_{D}$, HNSW+${}_{D}$, and TkQ+${}_{D}$, we set hyperparameter the number of objects retrieved (top-$k$) from 40,000, 20,000, 1,$0$, 5,000 to 2,000. This ensures the number of objects retrieved for reranking is consistent across different baselines, leading to a fair comparison. Other hyperparameter settings are fixed. We present the trade-off results on the NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 4 while other metrics and datasets have similar results. We have the following findings: (1) LIST outperforms other methods on both effectiveness and efficiency. This demonstrates LIST’s capability of maintaining effectiveness while accelerating querying speed. (2) The effectiveness of TkQ+${}_{D}$ is consistent as $k$ increases while time overhead increases significantly. This is due to the word mismatch issue discussed in Section 3.2. Many relevant objects without word overlap with the query remain hard to retrieve as $k$ increases. (3) HNSW+${}_{D}$ exhibits almost constant results. This is because HNSW builds a sparse proximity graph, limiting the reachable objects for each query. Therefore, when $k$ exceeds the number of reachable objects, its performance remains unchanged.
@@
Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$). We further investigate the impact of the number of clusters $cr$ to route queries and objects. We vary $cr$ from 1, 2, 3, to 4 for IVF+${}_{D}$ and our index, from 4, 5, 6, to 7 for IVFPQ+${}_{D}$. For other baselines, we set hyperparameter the number of objects retrieved (top-$k$) from 10,000, 20,000, 3,$0$, to 4,$0$ for a fair comparison like the previous trade-off study. Other hyperparameter settings are fixed. Similarly, we report the trade-off results on NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 5. According to the experiment results, LIST outperforms other methods in both effectiveness and efficiency under different $cr$ settings. Increasing $cr$ can enhance effectiveness, but it also leads to higher latency. In practice, users can adjust this parameter according to their needs.
Memory Consumption of LIST. The memory consumption of LIST is composed of three parts: the memory used by the proposed relevance model, the memory used by the proposed index, and the memory used for object text embeddings that are produced in advance. The experiment results are presented in Table 5, which demonstrates the remarkable memory efficiency of LIST. Compared with LSH+${}_{D}$, TkQ+${}_{D}$, and HNSW+${}_{D}$, LIST requires less memory. The reason is that our index stores only a lightweight MLP $c$-cluster classifier. This storage requirement is less than the memory consumption of the proximity graph of HNSW, hash tables of LSH, and the inverted file of the IR-Tree.
Table 5: Memory usage on three datasets (MB).
@@
5.4 Scalibility Study (RQ3)
Scalability Study. We further explore the scalability of LIST and the proposed relevance model LIST-R. We supplement the Geo-Glue dataset with new POIs crawled from Open Street Maps in Hangzhou. Since there is no ground-truth data for the relevance between the newly crawled POIs and the existing queries, we only report the efficiency of our proposed retriever LIST and our relevance model LIST-R. For the newly acquired POIs, we utilize the trained dual-encoder module and index to partition them into distinct clusters. We report the runtime of LIST and LIST-R on Figure 6. We observe that, as the number of objects increases, the runtime of LIST and LIST-R scales linearly.